evaluate the following integral
∫ 1 / x(3 + log x) dx

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello student,
Please find the answer to your question below

I = \int \frac{1}{x(logx+3)}dx
logx+3 = t
\frac{1}{x}dx = dt
I = \int \frac{1}{t}dt
I = log(t)+c
I = log(logx+3)+c